Electrons,each with charge e,move with speed v along a metal wire.the electric current in the wire is I.Plane P is perpendicular to the length of  wire.How many electrons pass through plane P in each second?

Grade:12th pass

2 Answers

Eshan
5 years ago
Dear student,

Charge passing the plane in each second=I
Therefore, number of electrons passing the plane in each second=\dfrac{I}{e}
yathartha gupta
5 years ago
Brother,
               Charge carriers are electrons  
      The equation represents that,
                  q=it and q=ne
              Thus, ne=it
 So,.        n= it/e
              According the question time t=1
        Hence,         number of electrons passing through out the plane each second = I/e